OpenStudy (anonymous):

slope intercept form of through (3,3) parallel to y=1/4x+4

OpenStudy (anonymous):

if the line is parallel to y = 1/4 x+4, then it has the same slope. so the new line is y = 1/4 x +c , with (x,y) = (3,3)